OpenStudy (e.cociuba):

Solve using elimination, graphing, or substitution: 4x - 3y= 1 y + 2x= 3

OpenStudy (nubeer):

re arrange your second equation .. it would become. y= 3-2x now put this value of y in your first equation and solve it...

OpenStudy (anonymous):

elimination: 1. re-write y+2x=3 2. choose either x or y to eliminate 3. when eliminating one they have to be equal but opposite ex: 2x, -2x 4. so lets say ur eliminating the x and u make them equal a shown above. you then do the math for "y" (im not done yet i also have to do the math lol)

OpenStudy (anonymous):

5. after finding y you repeat to find x.

OpenStudy (e.cociuba):

Ok I'm lost. What's with all the rearranging? :)

OpenStudy (anonymous):

ok , y+2x = 3 isnt in the same form as as 4x-3y=1, so you have to change it to the same form as 4x-3y=1

OpenStudy (e.cociuba):

Ok so it's 2x + y=3 4x - 3y=1 ??

OpenStudy (anonymous):

yeah, and after u get that u can choose either the x or y to eliminate

OpenStudy (luigi0210):

Graphing was an option so.. here you go ^_^' It has color

OpenStudy (anonymous):

You can rearrange either equation but rearranging y+2x=3 for y is the easiest. y=3-2x you put that value for y in the first equation 4x-3y=1 This gives 4x-3(3-2x)=1 4x-9+6x=1 4x+6x=10 10x=10 x=1 put x=1 into the second equation y=3-2(1) which gives you y=1 now you can test if this is correct by putting the values of x and y into the first equation 4(1)-3(1)=1

OpenStudy (anonymous):

@WilliamF oh ur doing substitution , right?

OpenStudy (anonymous):

@HELPKP Yeah

OpenStudy (anonymous):

@e.cociuba substitution never stuck to me,i liked elimination but at least you has all three ways to solve the equation

OpenStudy (anonymous):

i meant 'have'

OpenStudy (e.cociuba):

Ok so which should I do than lol?

OpenStudy (anonymous):

either one u do there're all right

OpenStudy (anonymous):

don't ever call me a liar again

OpenStudy (e.cociuba):

Well then y dont u explain y ur name went from Eric to Dylan suddenly.... Hmm?

OpenStudy (anonymous):

i was adopied out of foster care

OpenStudy (anonymous):

@e.cociuba Lairs will be lairs ;) #Trust NO ONE! xD

OpenStudy (e.cociuba):

True tht! :) @Cutie<3